In Feb 2006 in the Valley of the Kings in Egypt, a team of archaeologists uncovered the first tomb since King Tut's tomb was found in 1922. The tomb contained five wooden sarcophagi that contained mummies. The archaeologists believed that the mummies are from the 18th Dynasty, about 3300 to 3500 years ago. Determine the amount of carbon-14 that the mummies have lost. (Recall the half-life of carbon-14 is 5750 years.)
Answer by ankor@dixie-net.com(22740)

In Feb 2006 in the Valley of the Kings in Egypt, a team of archaeologists
uncovered the first tomb since King Tut's tomb was found in 1922.
The tomb contained five wooden sarcophagi that contained mummies.
The archaeologists believed that the mummies are from the 18th Dynasty, about 3300 to 3500 years ago.
Determine the amount of carbon-14 that the mummies have lost.
(Recall the half-life of carbon-14 is 5750 years.)
:
The carbon 14, formula: A = Ao*2^(-t/5750)
where
A = amt after t yrs (We will use per cent)
Ao = initial amt (let it = 1)
t = time in yrs
:
A = 1*2^(-3400/5750)
A = 2^-.5913
A = .6637 or 66.37% remain
therefore
100 - 66.37 = 33.63% has been lost